OpenStudy (anonymous):

solve this system: x - y = 0 2x - 2y + z + 2w = 4 y + w = 0 2z + w = 5

OpenStudy (anonymous):

if x=-y=0 2x-2y=0 hence z+2w=4 and 2z+w=5 solve to get w and z y=-w and x=y you can get all the variables.

OpenStudy (anonymous):

z=2 w=1 y=-1 x=-1

OpenStudy (anonymous):

You can write z in terms of w. And, you can write w in terms of y. Then, substitute in second equation. Now, you will get two equations in x and y. Solve them to get x and y. Then w and z from value of y.

OpenStudy (anonymous):

err in the first line the first sentence was x-y=0,a typo sorry.

OpenStudy (anonymous):

thank you , but how you get z y=-w ?

OpenStudy (anonymous):

I did not mean that that z is a part of the previous line it is just y=-w :)

OpenStudy (anonymous):

haha , i'm sorry .. thank you so much